OpenStudy (anonymous):

WILL FAN AND MEDAL! Given ΔGHI with G(4, –3), H(–4, 2), and I(2, 4), find the perpendicular bisector of Segment HI in standard form.

OpenStudy (anonymous):

 slope of HI=\[\frac{ 4-2 }{ 2-(-4) }=\frac{ 2 }{ 6 }=\frac{ 1 }{ 3 }\] slope of perpendicular bisector\[=\frac{ -1 }{ \frac{ 1 }{ 3 } }=-3\] eq. of perpendicular bisector is y-3=-3(x+1) y-3=-3x-3 y=-3x